40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 9 11:46:01 2021, Anchal, Paco, Summary, AUX, Xend Green Laser PDH OLTF measurement image-6f2923a3-01ce-4d04-bc53-d8db0238e195.jpgimage-72223f4b-3b74-4574-a7ad-de6628a2c5e9.jpgX_Green_ARM_PDH_OLTF.pdf
    Reply  Thu Jun 10 14:01:36 2021, Anchal, Summary, AUX, Xend Green Laser PDH OLTF measurement loop algebra AUX_PDH_LOOP.pdf
       Reply  Mon Jun 14 18:57:49 2021, Anchal, Update, AUX, Xend is unbearably hot. Green laser is loosing lock in 10's of seconds XAUX_PDH_Err_In_ASD.pdfXAUX_PZT_Out_Mon_ASD.pdf
       Reply  Tue Jun 15 15:26:43 2021, Anchal, Paco, Summary, AUX, Xend Green Laser PDH OLTF measurement loop algebra, excitation at control point AuxPDHloop.pdf
          Reply  Fri Jun 18 10:07:23 2021, Anchal, Paco, Summary, AUX, Xend Green Laser PDH OLTF with coherence XEND_PDH_OLTF_with_Coherence.pdfBeta_Amp.pdf
Message ID: 16200     Entry time: Mon Jun 14 18:57:49 2021     In reply to: 16197
Author: Anchal 
Type: Update 
Category: AUX 
Subject: Xend is unbearably hot. Green laser is loosing lock in 10's of seconds 

Working in Xend with mask on has become unbearable. It is very hot there and I would really like if we fix this issue.


Today, the Xend Green laser was just unable to hold lock for longer than 10's of seconds. The longest I could see it hold lock was for about 2 minutes. I couldn't find anything obviously wrong with it. Attached are noise spectrums of error and control points. The control point spectrum shows good matching with typical free running laser noise.

Are the few peaks above 10 kHz in error point spectrum worrysome? I need to think more about it in a cooler place to make sure.

I wanted to take a high frequency spectrum of error point to make sure that higher harmonics of 250 kHz modulation frequency are not leaking into the PDH box after demodulation. However, the lock could not be maintained long enough to take this final measurement. I'll try again tomorrow morning. It is generally cooler in the mornings.


This post is just an update on what's happening. I need to work more to get some meaningful inferences about this loop.

Attachment 1: XAUX_PDH_Err_In_ASD.pdf  59 kB  Uploaded Mon Jun 14 20:05:49 2021  | Hide | Hide all
XAUX_PDH_Err_In_ASD.pdf
Attachment 2: XAUX_PZT_Out_Mon_ASD.pdf  62 kB  Uploaded Mon Jun 14 20:06:08 2021  | Hide | Hide all
XAUX_PZT_Out_Mon_ASD.pdf
ELOG V3.1.3-